OpenStudy (anonymous):

find all real numbers x such that x^4 - 2x^2 - 15=0

OpenStudy (anonymous):

factor--> (x^2-5)(x^2+3)=0, set x^2-5=0, x^2+3=0, solve for each x.

OpenStudy (anonymous):

x = square root(5) and x = -square root(5)